Circular No. 6912

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ce IN MCG –4-24-19 S. Woodings, R. Martin, and A. Williams report their discovery of an

apparent supernova (R ∼ 16) on an image taken through thin cloud on May 19.50 UT with the Perth-Lowell 0.61-m reflector in the course of the Perth Astronomy Research Group’s automated supernova search (confirmed with an image taken on May 21.7). J. Biggs and A. Verveer provide the following position for SN 1998ce from unfiltered CCD images obtained around May 21.461 with the 0.25-m Mike Candy Telescope at Perth: α = 10^{h}10^{m}34^{s}.70 ± 0^{s}.02, δ = −25^{o}49′30′′.8 ± 0′′.3 (equinox 2000.0; mag 16.0), which is ∼ 9′′ west and 8′′ south of the center of MCG –4-24-19. The last image of the field was taken on May 3, with no object visible to mag ∼ 19.

GRB 980519 L. Nicastro, Istituto di Fisica Cosmica ed Applicazioni all’Informatica,

CNR, Palermo; L. A. Antonelli, BeppoSAX Science Data Center, Rome; G. Celidonio, M. R. Daniele, C. De Libero, and G. Spoliti, BeppoSAX Science Operation Center, Rome; L. Piro, Istituto Astrofisica Spaziale, CNR, Rome; and E. Pian, Istituto di Tecnologie e Studio delle Radiazioni Extraterrestri, CNR, Bologna, on behalf of the BeppoSAX team, report: “The BeppoSAX Wide Field Camera error box of GRB 980519 (IAUC 6910) was observed with the BeppoSAX Narrow Field Instruments starting on May 19.920 UT (9.7 hr after the burst). A previously unknown x-ray source, 1SAX J2322.3+7716, was detected by the LECS and MECS instruments (units 2 and 3) at α = 23^{h}22^{m}22^{s}, δ = +77^{o}16′.1 (equinox 2000.0; error radius 50′′). This position is 1′.2 from the centroid of the WFC position and is compatible with the optical transient detected by Jaunsen et al. (GCN 78; α = 23^{h}22^{m}21^{s}.46, δ = +77^{o}15′43′′.0, equinox 2000.0; I ∼ 19.5 on May 19.87). The source exhibits a non-monotonic decay, steeper during the first 4–5 hr of the 1.2-day observation. The average count rate of the first 4 hr is 0.007 ± 0.001 counts/s (2–10 keV) in the two MECS units, corresponding to (3.8 ± 0.6) × 10^{−13} erg cm^{−2} s^{−1} (assuming a Crab-like energy spectrum). At the end of the observation, the source had faded by a factor of ∼ 4. These results strongly suggest that 1SAX J2322.3+7716 is the x-ray afterglow of GRB 980519.”

1998 May 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green